OpenStudy (anonymous):

find solution set to t−3t+6−−−−−√=−2

OpenStudy (anonymous):

\[t-\sqrt{3t+6} = -2\]

OpenStudy (anonymous):

you want to fin the value of t?

OpenStudy (anonymous):

yes both solutions

OpenStudy (anonymous):

ok... first transpose -square root of 3t + 6 to the right side, and transpose -2 to the left soit gives you a new equation that are easy to handle... its t + 2 = square root of 3t+6...

OpenStudy (tkhunny):

#1 Consider the Domain \(3t+6 \ge 0\) #2 Isolate the radical on one side. Why are we posting the same question twice?

OpenStudy (anonymous):

raise both sides of the equation to the power of to to ditch the radical sign...

OpenStudy (anonymous):

power of 2 rather

OpenStudy (anonymous):

can someone draw this out or type the equation so i can get a visual picture

OpenStudy (tkhunny):

Did you do Step #2? #2 Isolate the radical on one side. Do that, and you will have a visual picture.

OpenStudy (anonymous):

t+2 = \[\sqrt{3t+6}\]

OpenStudy (tkhunny):

@retep If you use parentheses () rather than brackets [], that won't happen. :-)

OpenStudy (anonymous):

@retep then I square both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you expand the left side (t+2)^2 ..it gives you t^2 + 4t + 4, and the right side, when you raise it in the 2nd power the square root will be eliminated and it gives you a 3t + 6 only... got already

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok i figured it out thank you

OpenStudy (anonymous):

YW